# From Prompt Injections to Protocol Exploits: Threats in LLM-Powered AI Agents Workflows

### v1

@@ VII-F Evolutionary Coding Agents and Emerging Security Challenges
Novikov et al. [70] introduce AlphaEvolve, an autonomous evolutionary coding agent designed to push the boundaries of Large Language Models (LLMs) in addressing complex scientific and engineering challenges. AlphaEvolve operates by orchestrating a pipeline of LLMs that iteratively modify and improve source code based on feedback from evaluators. This evolutionary process enables the agent to optimize algorithms and potentially uncover novel, high-performance solutions. The authors demonstrate AlphaEvolve’s broad applicability through its successful deployment on critical computational problems, including the optimization of data center scheduling algorithms, simplification of hardware accelerator circuits, and acceleration of LLM training. Impressively, AlphaEvolve also discovered provably correct algorithms that surpass existing state-of-the-art methods in both mathematics and computer science, underscoring its potential to automate scientific discovery at scale.
While AlphaEvolve represents a significant advancement in automated coding agents, its architecture also introduces new vulnerabilities and security risks that must be carefully considered. The autonomous modification of code, especially in critical systems, creates potential attack vectors such as introducing subtle backdoors, exploiting evaluator feedback loops, or undetected generation of unsafe code. Moreover, the iterative nature of AlphaEvolve could amplify vulnerabilities if malicious code modifications are reinforced over successive generations. Future research should prioritize the development of secure evaluation mechanisms, formal verification of generated code, and techniques for detecting adversarial manipulation within the evolutionary process. Promising directions also include implementing strict access controls, real-time monitoring, and integrating secure coding standards to ensure that autonomous coding agents, such as AlphaEvolve, contribute to scientific and engineering progress without compromising system integrity.
@@ VII-G Robustness and Scalability of Defense Mechanisms
Despite the proliferation of input sanitization and prompt-filtering techniques, adversaries continue to discover new injection and jailbreak strategies that evade existing defenses. Current solutions often rely on static rules or model‐specific heuristics, which break down when faced with adaptive attackers who leverage long contexts, compositional prompts, or multi‐modal perturbations. Furthermore, most defense evaluations rely on limited benchmarks that fail to capture the full spectrum of real-world inputs, resulting in gaps in our understanding of LLM vulnerabilities.
Future work could prioritize the development of scalable, modality‐agnostic defense frameworks that offer formal guarantees of robustness. This includes adversarial training regimes that integrate continuous red-teaming (e.g., jailbreak fuzzers) into the development lifecycle, formal verification methods to certify sanitizer and filter correctness, and cross-model transfer evaluation to ensure defenses generalize beyond a single LLM implementation. Finally, building open, extensible benchmarks encompassing text, code, image, and audio attacks will be critical for measuring progress and guiding the community toward truly resilient systems.
@@ VII-H Securing Model Parameters and Memory Integrity
Model compromise attacks—such as backdoors and data‐ or memory‐poisoning—pose a serious threat because they persist beyond individual queries and can be nearly impossible to detect at inference time. Stealthy triggers embedded during fine‐tuning or through external memory banks can lie dormant until activated, undermining trust in the model’s outputs. Moreover, dynamic encryption techniques (e.g., DemonAgent) and composite backdoors that scatter triggers across inputs make traditional audit and anomaly detection insufficient.
Addressing these challenges requires new tools for end‐to‐end pipeline auditing and runtime integrity verification. Promising directions include: (1) cryptographic attestation of training artifacts and prompt templates to ensure provenance; (2) differential privacy and noise injection in memory retrieval to limit an attacker’s ability to influence the agent’s stored context; (3) continuous monitoring of model internals using lightweight trapdoors or canary triggers that signal unauthorized backdoor activations; and (4) tooling for transparent inspection and rollback of memory banks in deployed agents. By combining these approaches, we can begin to build models and agents whose internal state is both auditable and tamper‐resistant.
VII-I Resilience in Multi-Agent and Federated Environments
@@ VIII Conclusion
Our comprehensive taxonomy and threat survey provide a foundation for designing robust defense mechanisms and establishing best practices for resilient, secure LLM-agent workflows. We hope this work will guide future research and practical development toward closing the identified security gaps and enabling trustworthy, autonomous AI agents in real-world deployments.

### v2

@@ VII-F Evolutionary Coding Agents and Emerging Security Challenges
Novikov et al. [novikov2025alphaevolve] introduce AlphaEvolve, an autonomous evolutionary coding agent designed to push the boundaries of Large Language Models (LLMs) in addressing complex scientific and engineering challenges. AlphaEvolve operates by orchestrating a pipeline of LLMs that iteratively refine source code based on evaluators’ feedback. This evolutionary process enables the agent to optimize algorithms and potentially uncover novel, high-performance solutions. The authors demonstrate AlphaEvolve’s broad applicability through its successful deployment on critical computational problems, including the optimization of data center scheduling algorithms, simplification of hardware accelerator circuits, and acceleration of LLM training. Impressively, AlphaEvolve also discovered provably correct algorithms that surpass existing state-of-the-art methods in both mathematics and computer science, underscoring its potential to automate scientific discovery at scale.
While AlphaEvolve represents a significant advancement in automated coding agents, its architecture also introduces new vulnerabilities and security risks that must be carefully considered. The autonomous modification of code, especially in critical systems, creates potential attack vectors, such as the introduction of subtle backdoors, the exploitation of evaluator feedback loops, or the undetected generation of unsafe code. Moreover, the iterative nature of AlphaEvolve could amplify vulnerabilities if malicious code modifications are reinforced over successive generations. Future research should prioritize the development of secure evaluation mechanisms, formal verification of generated code, and techniques for detecting adversarial manipulation within the evolutionary process. Promising directions also include implementing strict access controls, real-time monitoring, and integrating secure coding standards to ensure that autonomous coding agents, such as AlphaEvolve, contribute to scientific and engineering progress without compromising system integrity.
@@ VII-G Robustness and Scalability of Defense Mechanisms
Despite the proliferation of input sanitization and prompt-filtering techniques, adversaries continue to discover new injection and jailbreak strategies that evade existing defenses. Current solutions often rely on static rules or model‐specific heuristics, which break down when faced with adaptive attackers who leverage long contexts, compositional prompts, or multi‐modal perturbations. Furthermore, most defense evaluations rely on limited benchmarks that fail to capture the full range of real-world inputs, leaving gaps in our understanding of LLM vulnerabilities.
Future work could prioritize the development of scalable, modality‐agnostic defense frameworks that offer formal guarantees of robustness. This includes adversarial training regimes that integrate continuous red-teaming (e.g., jailbreak fuzzers) into the development lifecycle, formal verification methods to certify the correctness of sanitizers and filters, and cross-model transfer evaluation to ensure defenses generalize beyond a single LLM implementation. Finally, building open, extensible benchmarks encompassing text, code, image, and audio attacks will be critical for measuring progress and guiding the community toward truly resilient systems.
@@ VII-H Securing Model Parameters and Memory Integrity
Model-compromise attacks—such as backdoors and data- or memory-poisoning—pose a serious threat because they persist beyond individual queries and can be nearly impossible to detect at inference time. Stealthy triggers embedded during fine‐tuning or through external memory banks can lie dormant until activated, undermining trust in the model’s outputs. Moreover, dynamic encryption techniques (e.g., DemonAgent) and composite backdoors that scatter triggers across inputs make traditional audit and anomaly detection insufficient.
Addressing these challenges requires new tools for end‐to‐end pipeline auditing and runtime integrity verification. Promising directions include: (1) cryptographic attestation of training artifacts and prompt templates to ensure provenance; (2) differential privacy and noise injection in memory retrieval to limit an attacker’s ability to influence the agent’s stored context; (3) continuous monitoring of model internals using lightweight trapdoors or canary triggers that signal unauthorized backdoor activations; and (4) tooling for transparent inspection and rollback of memory banks in deployed agents. By combining these approaches, we can begin building models and agents whose internal states are both auditable and tamper-resistant.
VII-I Future Mitigation Directions for MCP SDK Security
Another critical open challenge concerns the secure design and deployment of Model Context Protocol (MCP) Software Development Kits (SDKs), which often serve as the backbone for agent-to-tool integrations. To enhance their robustness, developers should adopt mitigation-oriented engineering practices such as enforcing least-privilege permissions, using mutual authentication and encrypted communications, validating structured data schemas, monitoring runtime anomalies, and maintaining cryptographically verifiable provenance logs. Furthermore, sandboxing plugin execution and auditing dependencies helps prevent malicious package injection and ensures the long-term resilience of MCP-based ecosystems. These practical measures provide a roadmap for strengthening SDK security and lay the foundation for trustworthy, large-scale agentic infrastructures.
VII-J Resilience in Multi-Agent and Federated Environments
@@
The term resilient multi-agent environments refers to systems capable of maintaining functionality and coordination in the face of both accidental faults and deliberate adversarial disruptions. Specifically, resilience encompasses fault tolerance—ensuring continued operation despite node or communication failures—and adversarial robustness—preserving secure interactions even when agents are exposed to manipulation, poisoning, or collusion.
@@ VIII Conclusion
Our comprehensive taxonomy and threat survey provide a foundation for designing robust defense mechanisms and establishing best practices for resilient, secure LLM-agent workflows. This framework provides practical value for developers securing agent workflows, researchers designing defenses, and policymakers establishing safety standards for LLM-based systems. We hope this work will guide future research and practical development toward closing the identified security gaps and enabling trustworthy, autonomous AI agents in real-world deployments.